OpenStudy (anonymous):

Simplify this complex fraction; (18x-6)/(9x5)/(15x+5)/(21x2)

OpenStudy (anonymous):

\[(18x-6)/(9x^5)/(15x+5)/(21x^2)=6(3x-1)21x^2/5(9x^5)(3x+1)\]\[=2(3x-1)/15x^3(3x+1)\]